OpenStudy (anonymous):

element x decays into element y with a half life of 20 years. suppose that a sample of element x was found to contain 70% of element y. estimate the age of this sample.I don't know how to start this, Thanks!

OpenStudy (john_es):

I think it should be a good start to put the relation, \[N=N_0e^{-\lambda t}\]

OpenStudy (john_es):

Then, it should be interesting to see that there is still 30% of element x, so \[N=0.3N_0\]

OpenStudy (anonymous):

is this simply the decay formula? p=poe^-rt?

OpenStudy (john_es):

Yes, I would say it.

OpenStudy (anonymous):

ok. but this is a differential equations class.. i dont understand.

OpenStudy (john_es):

Oh, for sure. Differential equations are coneccted with the decay process, \[dN=-\lambda N dt\]

OpenStudy (anonymous):

what does the differentiation mean?

OpenStudy (john_es):

Also you can start from this formula, integrate and obtain the decay formula, and then solve.

OpenStudy (john_es):

This is the "differential", so this is an exact differential equation that can be integrated, easily, \[\frac{dN}{N}=-\lambda dt\Rightarrow \int_{N_0}^N\frac{dN}{N}=-\lambda \int_0^t dt\]

OpenStudy (anonymous):

aha. i understand that i think. i'm sorry would you mind going over this problem with me in steps?

OpenStudy (john_es):

Yes, when you integrate the last formula you obtain, \[\ln(N/N_0)=-\lambda t\Rightarrow N=N_0e^{-\lambda t}\]

OpenStudy (john_es):

Now you need the conditions of the problem.

OpenStudy (anonymous):

can i just start the solution with N=N0e^rt ?

OpenStudy (john_es):

I think yes.

OpenStudy (anonymous):

and the conditions would it be N0=.3 and r=20 yearS?

OpenStudy (john_es):

Better, \[N=0.3N_0\]And for the halflife \[r=\lambda=\frac{\ln2}{T}\]with T=20 years.

OpenStudy (john_es):

r or λ is the constant of desintegration, that is related to the halflife. You can demonstrate this relation easily. But I give you a link where you can read it quietly. http://www.math24.net/radioactive-decay.html

OpenStudy (anonymous):

thanks so much. i will read over this and hopefully i will be able to get an answer

OpenStudy (john_es):

Yes, you will be able to do it, and in case you have any more doubts, tell me.

OpenStudy (anonymous):

ok so i got lamda to equal (1/20)*ln 2 and the equation to end up looking like this: 0.3 =e^(1/2ln2)t is that right?

OpenStudy (john_es):

\[0.3 =e^{(ln2/20)t}\]

OpenStudy (john_es):

with the minus sign in the exponent I forgot.

OpenStudy (anonymous):

why is it ln 2/20?

OpenStudy (john_es):

Because, \[\lambda=\frac{\ln 2}{T}=\frac{\ln 2}{20}\]

OpenStudy (john_es):

I would then apply logarithms, \[0.3=e^{-\frac{\ln2}{20}t}\Rightarrow \ln(0.3)=-\frac{\ln2}{20}t\Rightarrow t=-\frac{20\ln(0.3)}{\ln 2}\]

OpenStudy (anonymous):

i got it thanks so much!